OpenStudy (anonymous):

What are the coordinates of the vertex of the graph of the function y = –3x² –12x + 3? (–2, 29) (–2, 15) (2, –9) (3, –6)

OpenStudy (anonymous):

x=-b/2a; -(-12)/(2*-3) = 12/-6 = -2. The x coordinate is -2. Now find out what the y value at -2 is by plugging in -2 into the equation. (Remember variables have implied parenthesis!) f(-2)=-3(-2)^2-12(-2)+3 = -3(4)-(-24)+3 = -12+24+3 = 15. 15 is the y coordinate. This gives us (-2, 15). Evidence: http://www.wolframalpha.com/input/?i=y+%3D+%E2%80%933x%C2%B2+%E2%80%9312x+%2B+3 (see parabola properties)